Christos scored 80% in his automotive exam. There were 25 questions.

a. how many questions did he get right?
b. How many questions did he get wrong?
1
Expert's answer
2012-12-10T10:50:55-0500
Christos scored 80% in his automotive exam. There were 25 questions.
&
a. how many questions did he get right?

As Christos scored 80% of all the questions, then he got right

R = 25 * (80%/100%) = 25 * 0.8 = 20

questions.

b. How many questions did he get wrong?

As Christos scored 80% of all the questions, then he failed 100%-80% = 20% of them and so he got wrong

R = 25 * (20%/100%) = 25 * 0.2 = 5

questions.